OpenStudy (anonymous):

When an object is removed from a furnace and placed in an environment with a constant temperature of 23°C, its core temperature is 720°C. The temperature of the object t hours after it has been removed from the furnace is described by the functionF (t) = 23 + 697e−0.261t Based on this function, how warm will the object be after five hours i got 195 but that cant be right not a choice

OpenStudy (anonymous):

unless it is none of the above just check plz thanks

OpenStudy (anonymous):

Please let me know if there are any exponents in that formula

OpenStudy (anonymous):

The reason no one has helped you so far is the formula doesn't seem clear

OpenStudy (anonymous):

e^

OpenStudy (anonymous):

Looking this over I'm assuming the equation is the following: \[F(t) = 23 + 697e ^{-0.261t}\] If so plugging in 5 hours you get the following: \[F(5) = 23 +697e ^{-0.261*5}= 212\]